AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 13 of 33  ELOG logo
ID Date Author Subject Text Attachments
  419   Fri Jul 14 08:27:17 2023 TDCARME open - inspecting seal
09.20 CARME chamber at atmospheric pressure
as expected.
  
  418   Thu Jul 13 08:46:44 2023 TDReplacing wire seal
09.30 CARME chamber status
 20230713_135351.jpg20230713_135416.jpg 
  417   Thu Jul 13 00:12:45 2023 JM, CBCARME closedThe CARME chamber was closed and pushed
back into position on the ring. 

Prior to moving back onto the ring,
 Bias110723.pngfullyin.jpgfullyin-measured.jpg 
  416   Thu Jul 13 00:12:19 2023 JM, CBBellow replacedThe left (inside ring) bellow for the motors
rod was replaced.

Replacing the bellow involved;
 20230705_164523.jpg 
  415   Thu Jul 6 15:26:39 2023 JMLeak updateThe leak on the actuator arm bellows has
been attempted to be sealed using UHV vacuum
seal spray. The spray had the opposite to
 pumping.png 
  414   Wed Jul 5 19:27:13 2023 JMPump down and leak testCARME and the interaction chamber have
been pumping for ~26 hours, P=4.15e-7 mbar,
attached figures show the current pump down
 pumping_1.pngpumping-zoom.pngsignal-2023-07-05-15-13-30-512.jpg 
  413   Tue Jul 4 17:08:08 2023 JMLeak test and ventingThis morning the pressure achieved was
1.44E-7 mbar. Leak testing showed an overall
leak rate of 1.5E-10 mbar l/s, this is below
 20230704_135455.jpg20230704_174524.jpg 
  412   Mon Jun 26 14:49:05 2023 JM, TDCARME vented26/06/23

This morning the pressure was 4e-7
mbar. The bolts on the windows and the large
  
  411   Sun Jun 25 14:30:06 2023 JM, TDMotors bias testThe motors were moved using the pneumatic
only, in and out for several cycles to gauge
the effect on the vacuum. For the left arm
  
  410   Sun Jun 25 10:08:30 2023 JM, TDPump-down 23/06/23

Pumpdown started with the scroll
pump 17:15 . The turbo-pump was started 17:55.
 Figure_1.png20230625_120640.jpg20230625_120648.jpgpump-down.png 
  409   Sun Jun 25 09:41:06 2023 JM, TDNormal turbo-pump operationAttached are images showing the messages
for normal turbo-pump operation
 6x 
  408   Sat Jun 24 10:38:42 2023 JM, TDSection 9 closed, begin pumping 23/06/23

The CARME chamber and interaction
chamber were connected and section 9 has
 23-06-23-VI.png24-06-23-VI.png 
  407   Thu Jun 22 18:58:02 2023 JM, TDSlide back onto ringThe wire seal has been tightend further
to 160 Nm. Open neg feed-through flange has
also been fully tightend. 
 20230622_112534.jpg20230622_112827.jpg20230622_120307.jpg 
  406   Wed Jun 21 18:47:36 2023 JM, TDClosing CARMEInternal CARME installations complete. 


New strain reliefs (flange
side) installed on each cable harness.
2x
 20230621_104911.jpg20230621_104918.jpg20230621_110010.jpg21-06-23-VI.png 
  405   Fri May 26 09:08:06 2023 JM, CBDetector movement Hopefully we will not be opening the chamber
for quite some time so find attached images
of all four installed detectors and a video
 20230525_135756.jpgdetector-movement-in.mp4 
  404   Fri May 26 08:53:54 2023 JM, CBStrain reliefs (flange side) installedStrain reliefs on the flange side of the
cable harness have been installed for two
detectors (top right and bottom right detectors). 
 strain-relief-flange-side.jpgstrain-relief1_(1).mp420230526_140923.jpgV-I-plot.png 
  403   Wed May 24 09:54:56 2023 JM, CBMaximum detector positionsThe detectors were moved into the beam
axis using the motors to check the maximum
distance they can be moved in without collision
 9x 
  402   Wed May 24 08:35:12 2023 JM, CBStrain reliefs (detector side) installedThe new MACOR strain reliefs on the (detector
side) have been installed. See attached photos

The two pieces  of MACOR are
 20230524_092010.jpg20230524_092017.jpg20230524_092101.jpg20230524_093000.jpg20230524_092159.mp4 
  401   Mon May 15 07:34:22 2023 TDMonday 15 May08.35 DSSSD#3 bias -149.95V leakage current
-6.418uA temperature +19.6 deg C ( https://web-docs.gsi.de/~lestinsk/tempplot.php
)
  
  400   Sun May 14 11:55:12 2023 TDSunday 14 May12.44 DSSSD#3 bias -149.95V leakage current
-6.448uA temperature +19.7 deg C ( https://web-docs.gsi.de/~lestinsk/tempplot.php
)
  
ELOG V3.1.4-unknown